OpenStudy (anonymous):

Suppose that f is even and g is odd. a. Show that f o g is even. B. show that g o f is even .

zepdrix (zepdrix):

Hey Conad :) If f(x) is an even function, it satisfies this property: f(-x) = f(x). If g(x) is an odd function, it satisfies this property: g(-x) = -g(x). So what happens when you take their composition. f(g(x)) = ? Well again, let's plug in the negative x and see what happens to the whole thing: =f(g(-x)) Since g is an odd function, g(-x) will become -g(x). =f(-g(x)) Since f is an even function, f(-stuff) will become f(stuff). =f(g(x)) So we just showed that f(g(-x)) = f(-g(x)) = f(g(x)). The same property that even functions have.